Given :Assets-Jan.1,2015 $320 Liabilities- Jan. 1, 2015 $140 Assets-Dec. 31, 2015 $350 Revenues in 2015 $75 Expenses in 2015 $60 3. What are “ Liabilities” at December 31, 2015 if no withdrawals or investments were made during the year? (use the information from 2) A. $125 B. $110 C. $170 D. $155
Answer #1

Total assets=Total liabilities+Total equity

Total equity as on Jan.1,2015 =(320-140)

=$180

Addition to retained earnings=Revenue-Expenses

=(75-60)=$15

Hence equity as on December 31, 2015=Beginning equity+Addition to retained earnings

=(180+15)=$195

Hence liabilities at December 31, 2015 =350-195

=$155